I need help solving the following application problem:
Two added to the square root of the sum of a number and five is equal to six. Find the number?

Answer by lawdej(58)

2+square root(5+x)=6
square root (x+5)=6-2
square root (x+5)=4
x+5=16
x=16-5
x=11